During 2017, Whispering Winds Corp. reported net sales of $2550000, net income of $1270000, and depreciation expense of $120000. Whispering Winds also reported beginning total assets of $1130000, ending total assets of $1340000, plant assets of $1200000, and accumulated depreciation of $900000. Whispering Winds’s asset turnover (rounded) is

2.06 times.

1.90 times.

1.10 times.

1.10 times.
Answer #1

Average assets = (1130000+1340000/2) = 1235000

Asset turnover = Net sales/Average assets

= 2550000/1235000

Assets turnover = 2.06

So answer is a) 2.06 Times